OpenStudy (anonymous):

a superhero asked his nemesis how many cats she has. she answered with a riddle "Three-fourths of my cats plus seven" how many cats does the nemesis have? answer is 28 how do you get that answer

Directrix (directrix):

To get started, let C represent the superhero's total number of cats. "Three-fourths of my cats plus seven" indicates that the hero has the following number of cats: C = (3/4)C + 7 C - (3/4) C = 7 Simplify C - (3/4) C and set the result = 7. Solve for C. I think you'll find this to be correct when you come back online.: answer is 28 how do you get that answer @ineedbiohelpandquick